An article in Technometrics (1999, Vol. 41, pp. 202–211) studied the capability of a gauge by measuring the weight of paper. The data for repeated measurements of one sheet of paper are in the following table. Construct a 99% one-sided, upper confidence interval for the standard deviation of these measurements. Assume population is approximately normally distributed.

Round your answer to 3 decimal places.

Observations

3.481 3.448 3.485 3.475 3.472

3.477 3.472 3.464 3.472 3.470

3.470 3.470 3.477 3.473 3.474


σ ≤ _____________
1
Expert's answer
2020-07-20T18:51:51-0400

sample mean is x=(1/n)"\sum X_i" =52.08/15=3.472

The sample variance is given as,

s2=(1/(n-1))"\sum (X_i-x)^2=0.000966/14=0.000069"

confidence level, 1-"\alpha" =99%=0.99

degrees of freedom

df=14

"\chi^2=" 4.66

A 99% upper confidence bound for the variance is found as follows:

"\sigma^2<(n-1)s^2/\chi^2,"

"\sigma^2<(14)0.000069/4.66;\sigma^2<0.0002073;" "\sigma<0.0144"